OpenStudy (ayyookyndall):

P--> P'(–4, –7) for the glide reflection where the translation is (x, y) (x, y –4) and the line of reflection is x = –1. What are the coordinates of P? @StudyGurl14 @Data_LG2

OpenStudy (studygurl14):

Okay, first step. Find out how far point (-4,-7) is from x = -1

OpenStudy (ayyookyndall):

Okay, can you help me with that? I am still a little confused on it.

OpenStudy (studygurl14):

Okay, since it is x = -1, you use the x-value of the coordinate. You basically want to find the distance between x = -4 and x= -1

OpenStudy (studygurl14):

Can you do that?

OpenStudy (ayyookyndall):

Umm, I believe so. It would be -4, right?

OpenStudy (studygurl14):

Not quite... |-4 - (-1)| = |-4 + 1| = |-3| = ? (you have to take absolute value because there is no such thing as negative distance)

OpenStudy (ayyookyndall):

So it would be 3?

OpenStudy (studygurl14):

Yep. That means there are three units between points (-4, -7) and the reflection line x=-1. That also means that must be three units on the other side, between the reflection line and the reflected point. So that means double the distance (3) between the original point and the reflected point. So what we do is add 3(2) to the x-value of the coord. Can you do that? (this method may seem more involved, but it is good to know in case you don't have a graph handy or you don't want to draw it out)

OpenStudy (ayyookyndall):

Its 6

OpenStudy (studygurl14):

Right. So add 6 to just the x-coord of point (-4, -7)q

OpenStudy (studygurl14):

ignore the q

OpenStudy (ayyookyndall):

(2,-1)

OpenStudy (ayyookyndall):

If thats the answer, thats not one of the answers choices...

OpenStudy (studygurl14):

We're not done. We have to do the translation part now.

OpenStudy (ayyookyndall):

Ohh okay. How do we do that.

OpenStudy (studygurl14):

and it's not (2,-1) either

OpenStudy (studygurl14):

Add 6 to JUST the x-coord. So (2, -7)

OpenStudy (studygurl14):

The translation rule is y - 4. So, if you follow what happened last time...we add 4 to the y-coord.

OpenStudy (ayyookyndall):

So (2,-4)?

OpenStudy (studygurl14):

(2, -3) (-7 + 4 = -3)

OpenStudy (ayyookyndall):

I meant (2, -3) Opps

OpenStudy (studygurl14):

@Data_LG2 Does this look okay to you?

OpenStudy (anonymous):

yeah! :D purrty purrfect !

OpenStudy (ayyookyndall):

Okay, I have a few more questions. I will tag you guys! :)

OpenStudy (studygurl14):

:)